# Spatial distribution of organics in the Horsehead nebula: signposts of chemistry driven by atomic carbon

### v2

@@ 4.2 Chemistry of N-bearing species
In the case of HC3N, it is generally accepted that this molecule lacks efficient grain-surface formation pathways (Bergner et al. 2017, 2018). This may account for the absence of a filamentary emission structure at the PDR, in contrast to most of the species discussed in this work, which likely receive significant contributions from dust grain chemistry. In the gas phase, the formation of HC3N is primarily driven by hydrocarbon reactions such as C2H2 + CN and C2H + HCN $\rightarrow$ HC3N + H (Fukuzawa and Osamura 1997; Taniguchi et al. 2024). A less dominant, yet non-negligible, route involves C3H2 + N $\rightarrow$ HC3N (Murillo et al. 2022). In the Horsehead nebula, these hydrocarbon precursors (e.g., C2H, C2H2, and C3H2) are believed to originate mainly from gas-phase reactions initiated by C+ and CH, along with PAH photochemistry (Guzmán et al. 2015). However, an alternative origin linked to products of grain-surface C chemistry, such as CH4 (Lamberts et al. 2022; Tsuge et al. 2024), may also contribute (e.g., Hassel et al. 2008; Sakai and Yamamoto 2013; Loison et al. 2017).
@@
Figure 6: Schematic summary of the proposed scenario at the edge of the Horsehead nebula. The two crosses indicate the PDR and dense core positions, as in Fig. 2. At the PDR, the moderate FUV field enables the coexistence of C and CO, together with a sufficiently high $T_{\mathrm{dust}}$ to allow efficient diffusion of atomic C on grain surfaces. These conditions may favor the formation of H2CO, CH2CO, CH3CHO, HNCO, and CH3CN, likely through a combination of gas-phase and grain-surface pathways. Interestingly, for the two species with the lowest PDR abundances, CH3OH and HC3N, previous studies suggest ineffective photodesorption and inefficient grain-surface formation, respectively. In the case of the dense core, the FUV-shielded conditions limit CO photodissociation, and the low $T_{\mathrm{dust}}$ makes the diffusion of atomic C on grains inefficient. These conditions may favor the grain-surface formation of H2CO and CH3OH, and also contribute to the prevalence of HC3N in the gas phase, as it is less efficiently photodissociated.
@@ 4.3 Implications in other UV-irradiated environments
As a consequence of the higher FUV flux in the Orion Bar PDR ($G_{0}>10^{4}$, Marconi et al. 1998; Peeters et al. 2024), the degree of photodissociation is expected to be significantly greater than in the Horsehead PDR. Furthermore, dust grains in the Orion Bar are warmer ($T_{\mathrm{dust}}\simeq 40$$80$ K, Arab et al. 2012), which inhibits the formation of substantial ice mantles and thereby reduces the efficiency of grain surface chemistry. Indeed, we have found that most of the molecules analyzed in this work are less abundant in the gas phase of the Orion Bar PDR (Cuadrado et al. 2017) compared to the Horsehead PDR, with the remarkable exception of H2CO and CH3OH (see Table 4).
Given that CH3OH is predominantly formed on grains—with H2CO as one of its main precursors—a non-negligible grain-surface contribution deep inside the Orion Bar cannot be ruled out. Cuadrado et al. (2017) showed that pure gas-phase models fail to reproduce the observed abundances of H2CO and CH3OH, suggesting that the formation of organics on warm grains with limited ice mantles may still be feasible. Nevertheless, given the high abundance of C at the edge of the Orion Bar (Goicoechea et al. 2025) and the elevated dust temperatures, an efficient surface diffusion of atomic C on grains should play a significant role, but apparently is not the case. Hence, in contrast to the Horsehead PDR, where C can be sufficiently abundant on grains to compete with CO hydrogenation, the Orion Bar PDR appears to process ices dominated by CO rather than C, resembling those expected in the Horsehead dense core. This is, however, unlikely to occur in situ in the Bar owing to the high temperatures—though it may operate in dense cloud interiors, as evidenced by spatially resolved CH3OH observations (Leurini et al. 2010).
@@ 5 Summary and Conclusions
We have investigated the spatial distribution of simple and complex organic molecules at the FUV-irradiated edge of the Horsehead nebula. Using ALMA-7m interferometry combined with IRAM 30m single-dish data, we have mapped species such as C17O, HNCO, HC3N, CH3CN, CH2CO, and CH3CHO for the first time in this source, achieving an angular resolution of approximately $\sim$$15^{\prime\prime}$. Our analysis also incorporates deep single-dish observations of C17O and C18O from the WHISPER survey, along with PdBI maps of H2CO and CH3OH from Guzmán et al. (2013). Integrated line intensity profiles were extracted along the direction of the FUV radiation, and radiative transfer modeling was used to constrain the gas physical conditions and molecular column densities. The main results and conclusions are summarized as follows:
@@ Acknowledgements.
The authors thank the referee for the constructive comments that improved the content of this work. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2016.2.00027.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. Data reduction was carried out at the IRAM ARC node. C.H.-V. acknowledges support from the National Agency for Research and Development (ANID) – Scholarship Program through the Doctorado Nacional grant no. 2021-21212409. C.H.-V. and V.V.G. acknowledge support from the ANID – Millennium Science Initiative Program – Center Code NCN2024_001. V.V.G also gratefully acknowledges support from FONDECYT Regular 1221352, and ANID CATA-BASAL project FB210003. J.P. and M.G. acknowledge support by the French Agence Nationale de la Recherche through the DAOISM grant ANR-21-CE31-0010 and by the Thematic Action “Physique et Chimie du Milieu Interstellaire” (PCMI) of INSU Programme National “Astro”, with contributions from CNRS Physique & CNRS Chimie, CEA, and CNES. K.T.W. acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement no. 883867, project EXWINGS). J.R.G. thanks the Spanish MCINN for funding support under grant PID2023-146667NB-I00.
@@ Appendix A Observational details
Table 2 summarizes the observational details of the spectral setup used in the ALMA-7m observations.
Table 2: Spectral setup of the ALMA-7m observations [^10] Columns from left to right: baseband number, rest frequency (GHz) at the centre of the spectral window, targeted species, channel spacing (kHz), bandwidth (MHz), and the central frequency (GHz) of the TDM spectral window used in calibration.
@@ B.1 WHISPER CO isotopologues observations
Figure 7 presents the best-fit models accounting for the hyperfine structure of the C17O (top panels) and C18O (bottom panels) J=10J=1-0 and 212-1 rotational transitions, based on the WHISPER data analysis described in Sect. 3.2.1. The best-fit parameters and their statistical uncertainties for the PDR and dense core positions are shown in the middle of each panel. Averaging the results for the PDR and dense core yields $T_{\mathrm{ex}}(\mathrm{C^{17}O})\sim 15$ K and $N(\mathrm{C^{18}O})/N(\mathrm{C^{17}O})\sim 3.6$, which are the values adopted in Sect. 3.2.1 for the posterior analysis of the ALMA-7m+30m combined dataset.
@@
B.2 ALMA-7m+30m C17O observations
Figure 8 presents zeroth-moment maps of the hyperfine components of the C17O J=10J=1-0 rotational transition from the combined ALMA-7m+30m dataset. Only two maps are displayed, as the two hyperfine components at $112.359$ GHz are blended (see Table 3). The green and cyan squares mark the PDR and dense core positions, respectively, where the spectra were extracted for the calibration of the C17O-to-H2 conversion factor.
@@ B.2 ALMA-7m+30m C17O observations
Figure 8: Top row: ALMA-7m+30m zeroth-moment maps of the hyperfine components of the C17O J=10J=1-0 rotational transition resolved at the Horsehead edge. The rotation angle and red vertical line are the same as in Fig. 2. Contours are $5,10,20,30,50,100]\times\sigma$, where $\sigma$ is the zeroth-moment rms listed in Table 1. The dense core and PDR positions are marked with cyan and green squares, respectively, indicating the regions where the spectra were extracted. The beam size and a scale bar indicating $4000$ au are shown in the bottom left and bottom right corner, respectively, of each panel. Bottom row: C17O spectra extracted at the dense core and PDR positions, shown using matching colors to the map markers. The best-fit models are overlaid in red. The best-fit excitation conditions and associated uncertainties, presented in the top left corner of each panel, correspond to the 16th, 50th, and 84th percentiles of the posterior distributions.
B.3 ALMA-7m+30m CH3CN observations
@@
Figure 9: Integrated intensity profiles of the CH3CN J=65J=6-5 ($K=0,1,$ and $2$) transitions at the $\delta y$ positions of the dense core (left panel) and the PDR (right panel), as obtained from the ALMA-7m+30m observations. Each color corresponds to a different $K$-ladder transition, as indicated by the labels in the right panel. Observed data points are shown as circles, and the best-fit models are overlaid as solid lines in the corresponding colors. The red vertical line, the dotted vertical lines, and the horizontal gray bar are the same as in the left panels of Fig. 3.

### v3

@@ 4.2 Chemistry of N-bearing species
In the case of HC3N, it is generally accepted that this molecule lacks efficient grain-surface formation pathways (Bergner et al. 2017, 2018). This may account for the absence of a filamentary emission structure at the PDR, in contrast to most of the species discussed in this work, which likely receive significant contributions from dust grain chemistry. In the gas phase, the formation of HC3N is primarily driven by hydrocarbon reactions such as C2H2 + CN and C2H + HCN $\rightarrow$ HC3N + H (Fukuzawa and Osamura 1997; Taniguchi et al. 2024). A less dominant, yet nonnegligible, route involves C3H2 + N $\rightarrow$ HC3N (Murillo et al. 2022). In the Horsehead nebula, these hydrocarbon precursors (e.g., C2H, C2H2, and C3H2) are believed to originate mainly from gas-phase reactions initiated by C+ and CH, along with PAH photochemistry (Guzmán et al. 2015). However, an alternative origin linked to products of grain-surface C chemistry, such as CH4 (Lamberts et al. 2022; Tsuge et al. 2024), may also contribute (e.g., Hassel et al. 2008; Sakai and Yamamoto 2013; Loison et al. 2017).
@@
Figure 6: Schematic summary of the proposed scenario at the edge of the Horsehead nebula. The two crosses indicate the PDR and dense core positions, as in Fig. 2. At the PDR, the moderate FUV field enables the coexistence of C and CO, together with a sufficiently high $T_{\mathrm{dust}}$ to allow efficient diffusion of atomic C on grain surfaces. These conditions may favor the formation of H2CO, CH2CO, CH3CHO, HNCO, and CH3CN, likely through a combination of gas-phase and grain-surface pathways. Interestingly, for the two species with the lowest PDR abundances, CH3OH and HC3N, previous studies suggested ineffective photodesorption and inefficient grain-surface formation, respectively. In the case of the dense core, the FUV-shielded conditions limit CO photodissociation, and the low $T_{\mathrm{dust}}$ makes the diffusion of atomic C on grains inefficient. These conditions may favor the grain-surface formation of H2CO and CH3OH, and also contribute to the prevalence of HC3N in the gas phase as it is less efficiently photodissociated.
@@ 4.3 Implications in other UV-irradiated environments
As a consequence of the higher FUV flux in the Orion Bar PDR ($G_{0}>10^{4}$, Marconi et al. 1998; Peeters et al. 2024), the degree of photodissociation is expected to be significantly greater than in the Horsehead PDR. Furthermore, dust grains in the Orion Bar are warmer ($T_{\mathrm{dust}}\simeq 40$$80$ K, Arab et al. 2012), which inhibits the formation of substantial ice mantles and thereby reduces the efficiency of grain surface chemistry. We have found that most of the molecules analyzed in this work are less abundant in the gas phase of the Orion Bar PDR (Cuadrado et al. 2017) compared to the Horsehead PDR, with the remarkable exception of H2CO and CH3OH (see Table 4).
Given that CH3OH is predominantly formed on grains—with H2CO as one of its main precursors—a nonnegligible grain-surface contribution deep inside the Orion Bar cannot be ruled out. Cuadrado et al. (2017) showed that pure gas-phase models fail to reproduce the observed abundances of H2CO and CH3OH, suggesting that the formation of organics on warm grains with limited ice mantles may still be feasible. Nevertheless, given the high abundance of C at the edge of the Orion Bar (Goicoechea et al. 2025) and the elevated dust temperatures, an efficient surface diffusion of atomic C on grains should play a significant role, but apparently is not the case. Hence, in contrast to the Horsehead PDR, where C can be sufficiently abundant on grains to compete with CO hydrogenation, the Orion Bar PDR appears to process ices dominated by CO rather than C, resembling those expected in the Horsehead dense core. This is, however, unlikely to occur in situ in the Bar owing to the high temperatures—though it may operate in dense cloud interiors, as evidenced by spatially resolved CH3OH observations (Leurini et al. 2010).
@@ 5 Summary and Conclusions
We have investigated the spatial distribution of simple and complex organic molecules at the FUV-irradiated edge of the Horsehead nebula. Using ALMA 7m interferometry combined with IRAM 30m single-dish data, we have mapped species such as C17O, HNCO, HC3N, CH3CN, CH2CO, and CH3CHO for the first time in this source, achieving an angular resolution of approximately $\sim$$15^{\prime\prime}$. Our analysis also incorporates deep single-dish observations of C17O and C18O from the WHISPER survey, along with PdBI maps of H2CO and CH3OH from Guzmán et al. (2013). Integrated line intensity profiles were extracted along the direction of the FUV radiation, and radiative transfer modeling was used to constrain the gas physical conditions and molecular column densities. The main results and conclusions are summarized as follows:
@@ Acknowledgements.
The authors thank the referee for the constructive comments that improved the content of this work. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2016.2.00027.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. Data reduction was carried out at the IRAM ARC node. C.H.-V. acknowledges support from the National Agency for Research and Development (ANID) – Scholarship Program through the Doctorado Nacional grant no. 2021-21212409. C.H.-V. and V.V.G. acknowledge support from ANID – Millennium Science Initiative Program – Center Code NCN2024_001. V.V.G also gratefully acknowledges support from FONDECYT Regular 1221352, and ANID CATA-BASAL project FB210003. J.P. and M.G. acknowledge support by the French Agence Nationale de la Recherche through the DAOISM grant ANR-21-CE31-0010 and by the Thematic Action “Physique et Chimie du Milieu Interstellaire” (PCMI) of INSU Programme National “Astro”, with contributions from CNRS Physique & CNRS Chimie, CEA, and CNES. K.T.W. acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement no. 883867, project EXWINGS). J.R.G. thanks the Spanish MCINN for funding support under grant PID2023-146667NB-I00.
@@ Appendix A Observational details
Table 2 summarizes the observational details of the spectral setup used in the ALMA 7m observations.
Table 2: Spectral setup of the ALMA 7m observations [^10] Columns from left to right: Baseband number, rest frequency (GHz) at the center of the spectral window, targeted species, channel spacing (kHz), bandwidth (MHz), and the central frequency (GHz) of the TDM spectral window used in calibration.
@@ B.1 WHISPER CO isotopologues observations
Figure 7 presents the best-fit models accounting for the hyperfine structure of the C17O (top panels) and C18O (bottom panels) J=10J=1-0 and 212-1 rotational transitions, based on the WHISPER data analysis described in Sect. 3.2.1. The best-fit parameters and their statistical uncertainties for the PDR and dense core positions are shown in the middle of each panel. Averaging the results for the PDR and dense core yields $T_{\mathrm{ex}}(\mathrm{C^{17}O})\sim 15$ K and $N(\mathrm{C^{18}O})/N(\mathrm{C^{17}O})\sim 3.6$, which are the values adopted in Sect. 3.2.1 for the posterior analysis of the ALMA 7m+30m combined dataset.
@@
B.2 ALMA 7m+30m C17O observations
Figure 8 presents zeroth-moment maps of the hyperfine components of the C17O J=10J=1-0 rotational transition from the combined ALMA 7m+30m dataset. Only two maps are displayed, as the two hyperfine components at $112.359$ GHz are blended (see Table 3). The green and cyan squares mark the PDR and dense core positions, respectively, where the spectra were extracted for the calibration of the C17O-to-H2 conversion factor.
@@ B.2 ALMA 7m+30m C17O observations
Figure 8: Top row: ALMA 7m+30m zeroth-moment maps of the hyperfine components of the C17O J=10J=1-0 rotational transition resolved at the Horsehead edge. The rotation angle and red vertical line are the same as in Fig. 2. The contours are $5,10,20,30,50,100]\times\sigma$, where $\sigma$ is the zeroth-moment rms listed in Table 1. The dense core and PDR positions are marked with cyan and green squares, respectively, indicating the regions where the spectra were extracted. The beam size and a scale bar indicating $4000$ au are shown in the bottom left and bottom right corner, respectively, of each panel. Bottom row: C17O spectra extracted at the dense core and PDR positions, shown using matching colors to the map markers. The best-fit models are overlaid in red. The best-fit excitation conditions and associated uncertainties, presented in the top left corner of each panel, correspond to the 16th, 50th, and 84th percentiles of the posterior distributions.
B.3 ALMA 7m+30m CH3CN observations
@@
Figure 9: Integrated intensity profiles of the CH3CN J=65J=6-5 ($K=0,1,$ and $2$) transitions at the $\delta y$ positions of the dense core (left panel) and the PDR (right panel), as obtained from the ALMA 7m+30m observations. Each color corresponds to a different $K$-ladder transition, as indicated by the labels in the right panel. The observed data points are shown as circles, and the best-fit models are overlaid as solid lines in the corresponding colors. The red vertical line, the dotted vertical lines, and the horizontal gray bar are the same as in the left panels of Fig. 3.